OpenStudy (anonymous):

PLZ HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! The speed of a passenger train is 12 mph faster than the speed of a freight train. The passenger train travels 250 miles in the same time it takes the freight train to travel 190 miles. Find the speed of each train.

OpenStudy (anonymous):

times are equal .. time for passenger train is 250/(x+12) and freight is 190/x set 250/(x+12)=190/x and solve for x

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

250/(x+12)=190/x cross multiply to get 250x=190(x+12) 250x=190x + 190(12) 60x=190(12) etc

OpenStudy (anonymous):

x=38 so speeds are 38 and 50 mph

OpenStudy (anonymous):

witch is for witch?

OpenStudy (anonymous):

is 38 for the passenger train or freight train?